40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 12 15:14:09 2014, manasa, Update, General, X end AUX laser fiber setup ETMXtable.png
    Reply  Mon Dec 22 17:06:14 2014, manasa, Update, General, X end AUX laser fiber setup 
       Reply  Tue Dec 23 13:18:37 2014, manasa, Update, General, X end AUX laser fiber setup 
          Reply  Wed Jan 7 19:49:09 2015, manasa, Update, General, X end space issues ETMXspace.png
             Reply  Thu Jan 29 10:14:21 2015, manasa, Update, General, X end space issues photo.pngdrawing.png
                Reply  Wed Feb 4 21:09:43 2015, manasa, Update, General, X endtable work scheduled tomorrow 
                   Reply  Thu Feb 5 15:21:25 2015, manasa, Update, General, X endtable work  
                      Reply  Tue Feb 10 16:01:21 2015, manasa, Update, General, AUX X fiber coupled 72% AUXxTelescope.png
Message ID: 10996     Entry time: Tue Feb 10 16:01:21 2015     In reply to: 10981
Author: manasa 
Type: Update 
Category: General 
Subject: AUX X fiber coupled 72% 

Plan C finally worked. We have 1.454mW of AUX X light at the PSL table (2mW incident on the fiber coupler). 

Attached is the layout of the telescope.

What I did:

I stuck in Lens 1 (f=200mm) and measured the beam width after the focus of the lens at several points. I fit the data and calculated the beam waist and its position after this lens. 

I used the calculated waist and matched it with an appropriate lens and target (fiber coupler) distance. I calculated the maximum coupling efficiency to be 77% for Lens 2 with f=50mm and the fiber coupler placed at 20cm from the waist of Lens1. I was able to obtain 72% coupling after putting the telescope together.

I locked the arms, ran ASS and brought back GTRX to its usual optimum value of ~0.5 counts after closing. We also have the X arm beatnote on the spectrum analyzer.

Notes:

There are still a couple of things to fix. The rejected beam from the beam sampler has to be dumped using a razor blade.

 

ELOG V3.1.3-